40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Sun Nov 12 22:34:50 2017, Craig, awade, DailyProgress, NoiseBudget, Nov 12 Noisebudget PLL_OLG_SR785_Math_FMDevn_1kHz_PreampGain_2000_12-11-2017_195046_TF.pdfPLL_OLG_SR785_Math_FMDevn_1kHz_PreampGain_2000_12-11-2017_195046.tgz20171113_102854noiseBudget.pdf
    Reply  Mon Nov 13 15:45:19 2017, Craig, awade, HowTo, NoiseBudget, Nov 12 Noisebudget revised 20171113_151130noiseBudget.pdfPLLControlSignalSpectrum_Avg_30_FMDevn_1kHz_PreampGain_2000_Span_102p4kHz_12-11-2017_211211_Spectrum.pdf
       Reply  Mon Nov 13 18:06:32 2017, awade, Craig, DailyProgress, Schedule, Task list Nov 13, 2017 NorthLowerPeriscopeMirror_2017-11-13_16.48.15.jpgSouthLowerPeriscopeMirror_2017-11-13_16.49.35.jpg
          Reply  Mon Nov 13 21:53:25 2017, rana, DailyProgress, Schedule, Task list Nov 13, 2017 
Message ID: 1979     Entry time: Mon Nov 13 15:45:19 2017     In reply to: 1978     Reply to this: 1980
Author: Craig, awade 
Type: HowTo 
Category: NoiseBudget 
Subject: Nov 12 Noisebudget revised 

REVISED BOKEH NOV 12 2017 NOISEBUDGET.

I realized that the Nov 12 Spectrum 4 Hz frequency comb an strong spikes at 256 and 512 Hz are not true noise artifacts, but merely poor measurement post processing.

In particular, when spectra are stitched together, there is sometimes some overlap in the frequency domain.  When this happens, the data taken at lower resolution is supposed to be thrown out.  I neglected to do this, but now have fixed the code in iris to do this.  awade calls this the Hierarchical Chucker.  

When I measured a spectrum with high span up to 102.4 kHz, the first point is at 0 Hz, and the second is at 256 Hz, and so on.  These points are inflated with all low frequency noise, and should be thrown away if we have spectra taken at higher resolution.  When one includes these points, we get the blue 'Nov 12 2017 Beat' line below.  When the points are properly hierarchically chucked, you get the gold 'Nov 13 2017 Beat' line.  These are the same measurement, just with some different post processing, so they ought to lie right on top of one another, which they do except for the line artifacts.

I have also included a plot of all the spectra so people can see why a comb can appear where it shouldn't if you don't throw out the low frequency data of your low resolution measurements.  The '4 Hz comb' was actually just the spectra plotting script switching between the light orange and dark orange spectra in Plot 2.  The light orange spectrum is taken at 1 Hz resolution, while the dark orange is taken at 4 Hz. The light orange, at higher resolution, should be the only spectrum reported at low frequency.  But if you leave in the low frequency dark orange points, the higher noise floor emerges as a 4 Hz frequency comb.  This is why hierarchical chucking is a very important process.

Quote:

The lastest noisebudget is attached.  Here is the online Bokeh version.

We are reaching the levels that Evan and Tara reached in a small frequency region around 1 kHz.  We are supposed to do about 10 times better than them to reach the coating brownian noise floor for the AlGaAs coatings.

We took the PLL OLG via the SR785, as well as the transmission beatnote spectra.  I divided every spectrum by two because of the impedence matching effect.

The PLL was locked at 1 kHz/Vrms frequency modulation, and SR560 preamp gain of 2000.  This led to the PLL UGF of around 30 kHz.  With this high of a UGF we used the control signal to take the spectra, and set the Preamp Gain in calibratedBeatnoteSpectrum.py to one.

Tomorrow awade and I will set up the intensity stabilization servo to try and get even lower noise on transmission.  The error signal from the North cavity is breathing in an unusual way, and we hope to get rid of this excess noise.

 

Attachment 1: 20171113_151130noiseBudget.pdf  882 kB  | Hide | Hide all
20171113_151130noiseBudget.pdf
Attachment 2: PLLControlSignalSpectrum_Avg_30_FMDevn_1kHz_PreampGain_2000_Span_102p4kHz_12-11-2017_211211_Spectrum.pdf  134 kB  | Hide | Hide all
PLLControlSignalSpectrum_Avg_30_FMDevn_1kHz_PreampGain_2000_Span_102p4kHz_12-11-2017_211211_Spectrum.pdf
ELOG V3.1.3-